Suppose that shoe sizes of American women have a bell shaped distribution with a mean of 8.2 and a standard deviation of 1.49. using the empirical rule, what percentage of American women have shoe sizes that are less than 12.67? Please do not round your answer.
Answer #1

Solution Mean = 8.2 standard deviation = 1.49 P(x< 12.67)=P(X-4 < 12.67-8.21 1.49 =P(2<3) Plx< 12,67) = 0.0413 PIX<12-67)= 84thank you